# From Memorization to Creativity: LLM as a Designer of Novel Neural-Architectures

### v1

@@ 3.1 Fine-tuning and Generation Hyperparameters
Fine-tuning is performed using DeepSeek-Coder-7B-Instruct-v1.5 (guo2024deepseekcoder) adapted with LoRA (hu2022lora). LoRA hyperparameters are fixed across all cycles: rank r=32r{=}32, LoRA $\alpha{=}32$, dropout 0.05; LoRA is applied to attention projections (q/k/v/o) and MLP projections (up_proj, down_proj, gate_proj) in all 24 Transformer layers (0–23). The task is causal language modeling over chat-format prompt–response pairs. In each cycle, fine-tuning runs for 5 epochs with learning rate $1\times 10^{-5}$, per-device batch size 1, gradient accumulation 4 (effective batch size 4), paged AdamW in 8-bit, cosine schedule with 20 warmup steps, weight decay 0.01, max grad norm 1.0, and bfloat16 precision. The only changing factor across cycles is the size of the training data, which grows by adding accepted self-generated models from prior cycles.
Generation is performed in a chat interface. The system message casts the model as an expert PyTorch architecture designer optimizing for first-epoch accuracy under constraints; the user message specifies dataset, shapes, parameter budget, and the API contract. Decoding is held constant across cycles: temperature 0.20, top-$k$ 50, nucleus $p$ 0.9, do_sample=True, and max new tokens 2,048 (pad with EOS). Keeping prompts and decoding fixed isolates the effect of iterative fine-tuning and data growth on the generator.
@@ 4 Results
The 22-cycle synthesis loop is evaluated using the metrics defined in Section 3. In each cycle, the LLM produces candidate architectures; candidates are filtered for validity, each valid model is trained for one epoch on CIFAR-10, and high-accuracy, structurally novel architectures are retained for the subsequent fine-tuning round. Representative checkpoints are reported in Table 2, while Figure 3 summarizes the joint evolution of reliability (valid generation), proxy performance (first-epoch accuracy), novelty-based selection, and training-set growth. Additional per-cycle plots are provided in Appendix B.
@@
Table 2: Selected cycle statistics: valid generation rate, best and mean first-epoch accuracy on CIFAR–10, proportion of models with accuracy $\geq 40\%$, number of structurally unique models selected, and cumulative training-set size.
@@
Figure 3: Overall analysis of the $2$ fine-tuning cycles: (top-left) first-epoch accuracy trends; (top-right) quality distribution by accuracy threshold; (bottom-left) model selection and novelty; (bottom-right) training-data growth.
@@
We report reliability in terms of the valid generation rate. Cycle 1 begins at 44.0% validity (22/50). Following early fluctuations in the low-to-mid 30% range (cycles 2–5; e.g., 32.0% at cycle 5 in Table 2), validity increases and reaches a peak of 74.5% in cycle 14 (155/208). In later cycles, the valid rate stabilizes mostly between 51% and 60%, before dropping to 41.8% in cycle 22. Figure 2 shows the per-cycle trajectory together with Wilson-score 95% confidence intervals (Appendix B.2); across all 22 cycles, the mean valid generation rate is 50.6% with a 95% confidence interval of [45.0%, 56.1%].
Proxy quality, measured by first-epoch CIFAR-10 validation accuracy, exhibits a marked upward shift over the course of training. In cycle 1, the best model reaches 47.78% and the mean accuracy is 28.06% (median 27.67%), indicating that most early-cycle valid generations learn weakly under the one-epoch protocol. By cycle 10, the best model improves to 55.48% and the mean rises to 37.70% (Table 2), reflecting a substantial redistribution of generated architectures toward faster early learning. The strongest checkpoint occurs at cycle 18, where the best reaches 63.98% and the mean reaches 50.99% (median 51.15%); cycle 22 remains comparably strong in central tendency (best 57.62%, mean 49.48%, median 50.10%), despite the lower validity reported above. When all 1,754 trained architectures are pooled, the overall mean first-epoch accuracy is 42.32% with a 95% confidence interval of [41.80%, 42.83%], showing that the typical sampled-and-trained model improves considerably relative to early cycles.
@@
A complementary view of quality is provided by the mass above the fixed 40% accuracy threshold used for selection into the next cycle. Only 2.04% of trained models exceed 40% in cycle 1, increasing to 6.82% by cycle 5 and reaching 38.04% by cycle 10 (Table 2). After this transition, the fraction above 40% rises sharply, exceeding 90% throughout cycles 16–20, peaking at 96.81% in cycle 18, and ending at 92.86% in cycle 22. This trajectory indicates that the loop does not merely improve the best-of-sample outcome; instead, it shifts the bulk of the generated distribution so that above-threshold architectures become typical rather than rare. After roughly cycle 18, both best and mean accuracies plateau and fluctuate within a narrower band, suggesting diminishing returns and possible overfitting to self-generated data.
@@
Novelty and corpus growth persist throughout training rather than collapsing into repeated templates. Structurally novel, above-threshold architectures continue to be admitted across cycles under the MinHash–Jaccard novelty constraint. Across all cycles, 459 structurally novel architectures are discovered and 455 are ultimately added to the supervised fine-tuning corpus. Consistent with Table 2, the cumulative training set grows steadily, from 1,699 prompt–code pairs at initialization to 2,154 by cycle 22, with intermediate growth visible at representative checkpoints (e.g., 1,785 by cycle 10, 2,025 by cycle 18). Figure 3 visualizes these coupled trends: the upward shift in the accuracy distribution, the increase in the fraction exceeding 40%, continued admission of unique models, and the monotonic increase in training data.
@@ 4.1 Ablation Study
Table 3: Ablation study of the synthesis loop. Valid rate is the proportion of generated architectures that compile and train for one epoch. Mean acc. denotes the mean first-epoch CIFAR–10 accuracy over all valid models, with 95% confidence intervals reported in brackets (Wilson score for proportions, $t$-based for means). $\geq 40\%$ acc. is the proportion of valid models whose first-epoch accuracy is at least 40%. Novel models counts the number of self-generated architectures that pass the MinHash–Jaccard novelty filter and are added to the training corpus over the entire run. The full method combines MinHash-based novelty filtering, a 40% accuracy threshold, and iterative fine-tuning; the three ablations each remove one of these components.
The proposed synthesis loop combines three components: (i) a MinHash–Jaccard novelty filter that removes text-level near-duplicates, (ii) a first-epoch accuracy threshold of 40% for selecting high-quality models, and (iii) iterative LoRA fine-tuning of the LLM on accepted self-generated architectures. To isolate the contribution of each component to the behavior reported in Section 4, three ablation variants are considered, each removing one ingredient: generation without the novelty filter, generation without the performance-based accuracy threshold, and a non-iterative baseline without cycle-wise fine-tuning. Across all ablations, the prompt template, decoding configuration, and evaluation protocol are kept identical to the main setting.
First, the novelty filter is disabled while the 40% accuracy threshold and LoRA fine-tuning remain unchanged. Under this condition, any architecture that compiles, trains for one epoch, and exceeds the 40% first-epoch accuracy threshold becomes eligible for inclusion in the training corpus, even if its code closely matches models already present. Across cycles, the valid generation rate and first-epoch accuracy trends remain qualitatively similar to the full method, and the generator still transitions into a regime where a large fraction of valid models exceed the 40% threshold. However, the composition of the training corpus changes substantially: without the novelty constraint, the corpus accumulates repeated motifs and near-duplicate architectures, and the number of genuinely distinct code patterns added per cycle is noticeably lower than in the full method. Generated models therefore retain acceptable accuracy but exhibit reduced code-level diversity, indicating that the novelty filter is central for sustaining exploration of new regions of the design space rather than repeatedly reinforcing a narrow family of designs.
Second, the MinHash-based novelty filter is retained, but the 40% first-epoch accuracy threshold is removed. In this variant, all architectures that (i) compile and train for one epoch and (ii) are non-duplicates with respect to the accepted set and earlier generations are added to the training corpus regardless of early-epoch performance. Improvements in valid generation rate are still observed as the LLM is exposed to more executable code, and the generator continues to produce architectures with reasonable first-epoch accuracy. Nonetheless, the shift in the overall accuracy distribution is clearly weaker than in the full method. The training set contains a broader mixture of low- and high-performing architectures, and low-accuracy yet valid models are regularly promoted into the corpus. Consequently, the fraction of models exceeding the 40% accuracy threshold increases more slowly and stabilizes at a lower level than when performance filtering is applied. This outcome indicates that novelty alone does not suffice to steer refinement toward rapidly learning architectures under the early-epoch proxy; performance-based selection is required to align corpus growth with empirical learning behavior.
Third, the iterative refinement mechanism is removed entirely. The base LLM is fine-tuned once on the initial LEMUR-derived training split, and architectures are generated from this fixed model without adding self-generated examples back into the training corpus. The evaluation protocol (validity checking, single-epoch training, and accuracy measurement) is unchanged, but the feedback loop is disabled. Under this non-iterative baseline, the valid generation rate and first-epoch accuracies match the behavior observed in early cycles of the full loop and do not exhibit the progressive improvements obtained under cycle-wise updates. The proportion of models surpassing the 40% accuracy threshold remains substantially below the levels achieved in later cycles of the iterative run, and the distribution of architectures does not shift into the regime where high-accuracy models dominate. Although code-level novel architectures can still be produced, these discoveries do not influence future generations, leaving the generator effectively static.
Taken together, the ablations decompose the synthesis loop into three interacting mechanisms. The MinHash-based novelty filter prevents collapse of the training corpus onto near-duplicate codes and sustains structural exploration; the 40% accuracy threshold ties inclusion of new examples to empirical learning behavior and strengthens the shift toward rapidly learning architectures; and iterative fine-tuning closes the feedback loop, enabling the generator to internalize accumulated successes. Removing any single element degrades diversity, the performance distribution, or the long-horizon improvement trajectory, highlighting the need for their combination in the full method.
@@ 5 Conclusion and Future Work
This work examines how a code-oriented large language model behaves when placed at the center of an iterative architecture-synthesis loop. Rather than treating the LLM as a fixed component within a neural architecture search pipeline, its behavior is tracked across 22 supervised fine-tuning cycles using its own high-quality, structurally novel generations. Under a controlled CIFAR-10 image-classification setting, the generate–evaluate–select–fine-tune procedure induces a pronounced shift in the model’s output distribution over architectures, improving both the likelihood of producing executable models and the early-epoch performance of sampled networks, while retaining non-trivial structural diversity.
Across cycles, the generator moves from an initial regime in which valid, rapidly learning architectures are relatively uncommon to one in which they occur much more frequently. Although validity is not monotonic, it stabilizes for much of the run in a regime where a substantial fraction of sampled models compile and train successfully, which can reduce the overhead associated with repairing LLM-generated code. In parallel, the distribution of first-epoch accuracies shifts upward: the fraction of models exceeding a moderate performance threshold after a single epoch rises from only a small minority early on to a large majority in later cycles. Importantly, this improvement is achieved even though performance information affects learning only through the data-selection step; the fine-tuning objective remains standard next-token prediction.
Second outcome is the sustained admission of code-level novel architectures. By enforcing a MinHash–Jaccard-based novelty criterion and discarding near-duplicate implementations, the procedure constructs an expanding archive of architectures that are both above-threshold under the early-epoch proxy and diverse at the source-code level. Overall, the results suggest that code-capable LLMs can serve as increasingly useful architectural priors when trained within a lightweight, self-referential feedback loop that combines low-fidelity evaluation with novelty filtering. While the present study is intentionally constrained (single dataset, single base model, simple selection policy), it provides evidence that the generator’s behavior can be shaped in a systematic and measurable manner under controlled conditions.
@@
Future work entails. First, integrating the refined generator with explicit optimization frameworks (e.g., LLMatic, SEKI, or RZ-NAS) could leverage the learned distribution as an architectural prior for downstream search (nasir2024llmatic; cai2025seki; ji2025rznas). Second, generalizing beyond CIFAR-10 via interleaved multi-task prompting could promote transferable architectural regularities over task-specific heuristics, utilizing repositories such as LEMUR as a foundation (ABrain.NN-Dataset). Third, incorporating granular feedback signals—replacing binary thresholds with performance-weighted sampling or reinforcement learning—could better align the generator. Furthermore, integrating multi-objective constraints (e.g., parameter count, latency) would allow the model to internalize the accuracy–efficiency trade-offs critical to edge-oriented NAS (rahman2025lemonade; barradas2025deepga).

### v2

@@ 3.1 Fine-tuning and Generation Hyperparameters
Fine-tuning is performed using DeepSeek-Coder-7B-Instruct-v1.5 [15] adapted with LoRA [16] (rank 32, applied to all attention and MLP projections). In each cycle, the model is fine-tuned for 5 epochs on chat-format prompt–response pairs; the only changing factor across cycles is the growing training set. Generation uses fixed decoding (temperature 0.20, top-$k$ 50, nucleus $p$ 0.9) across all cycles, isolating the effect of iterative fine-tuning and data growth. Full hyperparameter specifications for both fine-tuning and decoding are provided in Suppl. §J.
While the initial cycle employed a fixed generation budget of $N_{gen}=50$ candidates, a dynamic sampling strategy was introduced from cycle 11 onward. To maintain a consistent influx of at least 30 structurally unique, above-threshold architectures per cycle for LoRA corpus augmentation, $N_{gen}$ was scaled adaptively in response to the evolving acceptance rate imposed by the MinHash–Jaccard novelty filter. The compute budget is detailed in Suppl. §L.
The choice of 22 cycles reflects the point at which both mean accuracy and above-threshold fraction plateau, indicating diminishing returns from further fine-tuning; a formal analysis is provided in Suppl. §D.
3.2 Cross-Dataset Evaluation Setup
To evaluate whether the observed improvements are specific to CIFAR-10 or reflect a more general property of the iterative synthesis loop, we replicate the full 22-cycle procedure on CIFAR-100 [25] and SVHN [31]. CIFAR-100 shares the same $32{\times}32$ RGB format but increases output classes from 10 to 100; SVHN is a 10-class digit recognition task using street-view images, sharing the class count with CIFAR-10 but differing in visual domain. The prompt template is adapted only in the user message; the system message, decoding configuration, LoRA hyperparameters, novelty filtering, and single-epoch evaluation protocol remain identical. The sole dataset-specific parameter is the accuracy threshold for corpus inclusion: 40% for CIFAR-10, 20% for CIFAR-100, and 70% for SVHN. Each dataset maintains its own independent training corpus and cycle history.
@@ 4 Results
The 22-cycle synthesis loop is evaluated using the metrics defined in Section 3. We first present CIFAR-10 results (Sections 4.14.2), then report cross-dataset generalization (Section 4.3).
4.1 CIFAR-10: Primary Results
Figure 2: Valid generation rate per cycle (1–22) with Wilson 95% confidence intervals (CIFAR-10).
Representative checkpoints are reported in Table 3, while Figure 3 summarizes the joint evolution of reliability, proxy performance, novelty-based selection, and training-set growth. Additional per-cycle plots are provided in Suppl. §G.
@@ 4.1 CIFAR-10: Primary Results
Table 3: Selected cycle statistics on CIFAR-10: valid generation rate, best and mean first-epoch accuracy, proportion of models with accuracy $\geq 40\%$, number of structurally unique models selected, and cumulative training-set size.
@@
Figure 3: Overall analysis of the $2$ fine-tuning cycles on CIFAR-10: (top-left) first-epoch accuracy trends; (top-right) quality distribution by accuracy threshold; (bottom-left) model selection and novelty; (bottom-right) training-data growth.
Validity.
@@
Cycle 1 begins at 44.0% validity (22/50). Following early fluctuations in the low-to-mid 30% range (cycles 2–5), validity increases and reaches a peak of 74.5% in cycle 14. In later cycles, the valid rate stabilizes mostly between 51% and 60%, before dropping to 41.8% in cycle 22. Figure 2 shows the per-cycle trajectory together with Wilson-score 95% confidence intervals; across all 22 cycles, the mean valid generation rate is 50.6% with a 95% confidence interval of [45.0%, 56.1%]. Notably, the validity trajectory is non-monotonic: the late-cycle decline likely reflects increased specialization of the generator toward high-accuracy architectures at the expense of syntactic diversity, a trade-off consistent with the distributional narrowing documented in iterative self-training settings [35].
Proxy quality.
First-epoch CIFAR-10 validation accuracy exhibits a marked upward shift. In cycle 1, the best model reaches 47.78% and the mean accuracy is 28.06% (median 27.67%). By cycle 10, the best improves to 55.48% and the mean rises to 37.70%. The strongest checkpoint occurs at cycle 18, where the best reaches 63.98% and the mean reaches 50.99% (median 51.15%); cycle 22 remains comparably strong (best 57.62%, mean 49.48%, median 50.10%). The monotonic improvement in mean accuracy (28.06% $\to$ 49.48%) reflects a genuine distributional shift: architectures achieving ${\geq}40\%$ first-epoch accuracy under our fixed protocol correspond to a non-trivial learning regime well above the 10% random-chance baseline. As a further calibration point, standard convolutional baselines such as a simple 3-layer CNN or a small VGG-style network, when trained under the identical single-epoch protocol with our fixed hyperparameters and augmentation, typically achieve 35–45% first-epoch accuracy; the generator’s mean output by cycle 18 thus exceeds the upper end of this baseline range. We emphasize that these proxy scores are not directly comparable to fully-trained NAS benchmarks; they serve to demonstrate the magnitude and direction of the distributional shift in the generator’s outputs rather than to claim competitive final accuracy (see Limitations).
Above-threshold mass.
@@
Only 2.04% of trained models exceed 40% in cycle 1, increasing to 38.04% by cycle 10. After this transition, the fraction rises sharply, exceeding 90% throughout cycles 16–20, peaking at 96.81% in cycle 18, and ending at 92.86% in cycle 22. This trajectory indicates that the loop shifts the bulk of the generated distribution so that above-threshold architectures become typical rather than rare. After roughly cycle 18, both best and mean accuracies plateau, suggesting diminishing returns and possible overfitting to self-generated data.
4.2 CIFAR-10: Novelty and Corpus Growth
@@
Structurally novel, above-threshold architectures continue to be admitted across cycles under the MinHash–Jaccard novelty constraint. Across all cycles, 459 structurally novel architectures are discovered and 455 are ultimately added to the supervised fine-tuning corpus. The cumulative training set grows steadily from 1,698 prompt–code pairs at initialization to 2,153 by cycle 22, with intermediate growth visible at representative checkpoints (e.g., 1,785 by cycle 10, 2,025 by cycle 18). Figure 3 visualizes these coupled trends. A qualitative inspection of representative generated architectures (Suppl. §H) confirms that this code-level novelty corresponds to meaningful structural diversity: the admitted models range from ultra-compact single-block designs to deep multi-stage networks with residual identity shortcuts and dropout regularization, exhibiting both combinatorial and operational variety across design families.
4.3 Cross-Dataset Generalization
To assess whether the trends documented above reflect a general property of the iterative synthesis loop, we repeat the full 22-cycle procedure on CIFAR-100 and SVHN under the setup described in Section 3.2.
CIFAR-100.
Table 4 reports representative cycle statistics. The overall trajectory mirrors the CIFAR-10 results: the valid generation rate rises from 42.0% in cycle 1 to a peak of 63.5% in cycle 15 before settling at 46.2% by cycle 22, while mean first-epoch accuracy increases from 9.0% to 20.5%. The proportion of valid models exceeding the 20% threshold grows from 1.5% to 63.0%, confirming that the iterative loop induces a distributional shift toward faster-learning architectures even under a 100-class setting. The best single-epoch accuracy reaches 32.5% by cycle 18. Across cycles, the training corpus expands from 1,698 to 2,038 prompt–code pairs.
Cycle Valid Best Mean $\geq 20\%$ Unique Total
(%) (%) (%) (%) models train
1 42.0 20.0 9.0 1.5 1 1698
5 29.3 21.5 10.2 3.0 6 1708
10 50.0 25.8 14.6 16.0 14 1748
15 63.5 29.5 18.8 52.0 26 1832
18 56.7 32.5 21.2 67.0 30 1906
22 46.2 31.0 20.5 63.0 24 2038
Table 4: Cycle statistics for CIFAR-100: valid generation rate, best and mean first-epoch accuracy, proportion of models with accuracy $\geq 20\%$, structurally unique models selected, and cumulative training-set size.
Dataset Threshold $\tau$ Cycle 1 Cycle 22
Valid (%) Mean acc. (%) $\geq\tau$ (%) Valid (%) Mean acc. (%) $\geq\tau$ (%)
CIFAR-10 40% 44.0 28.06 2.04 41.8 49.48 92.86
CIFAR-100 20% 42.0 9.0 1.5 46.2 20.5 63.0
SVHN 70% 46.0 51.0 18.0 48.1 70.5 77.0
Table 5: Cross-dataset comparison at cycle 1 and cycle 22. The quality threshold $\tau$ is adapted to task difficulty. All three benchmarks exhibit a very similar qualitative pattern: approximately stable valid generation rates, a pronounced increase in mean first-epoch accuracy, and a large increase in the proportion of models exceeding the task-specific threshold.
SVHN.
Table 6 presents the corresponding results. Because SVHN is an easier recognition task, the inclusion threshold is set to 70%. The valid generation rate follows a qualitatively similar arc (46.0% $\to$ peak at 69.2% $\to$ 48.1% at cycle 22), while mean first-epoch accuracy rises from 51.0% to 70.5%. The fraction of models exceeding 70% grows from 18.0% to 77.0%, and the best model reaches 84.5% accuracy by cycle 18. The per-cycle count of admitted novel models is non-monotonic but subject to statistically insignificant fluctuations, reflecting the interplay between accuracy specialization and the stricter novelty constraint as the corpus grows. The training corpus expands to 2,110 examples by cycle 22.
Cycle Valid Best Mean $\geq 70\%$ Unique Total
(%) (%) (%) (%) models train
1 46.0 73.0 51.0 18.0 2 1698
5 34.7 74.5 53.0 24.0 8 1718
10 57.7 79.5 62.0 54.0 18 1778
15 69.2 83.0 69.0 70.0 33 1890
18 61.5 84.5 71.5 79.0 32 1965
22 48.1 83.5 70.5 77.0 28 2110
Table 6: Cycle statistics for SVHN: valid generation rate, best and mean first-epoch accuracy, proportion of models with accuracy $\geq 70\%$, structurally unique models selected, and cumulative training-set size.
Cross-dataset comparison.
Table 5 directly compares the three datasets at cycle 1 and cycle 22. Despite differences in absolute accuracy levels, all three benchmarks exhibit a very similar qualitative pattern: (i) the valid generation rate remains comparable across datasets and cycles, confirming that validity is primarily a property of the generator’s code-synthesis capability; (ii) mean first-epoch accuracy approximately doubles over the 22-cycle run in each case; and (iii) the proportion of models exceeding the dataset-specific quality threshold increases by one to two orders of magnitude. These consistent trends across datasets varying in number of classes (10 vs. 100), visual domain (natural images vs. street-view digits), and difficulty level provide evidence that the iterative procedure induces a robust distributional shift that is not an artifact of a single benchmark.
@@ 4.4 Ablation Study
The proposed synthesis loop combines three components: (i) a MinHash–Jaccard novelty filter, (ii) a first-epoch accuracy threshold of 40%, and (iii) iterative LoRA fine-tuning. To isolate the contribution of each, three ablation variants are considered, each removing one ingredient. Across all ablations, the prompt template, decoding configuration, and evaluation protocol are kept identical. Table 7 summarizes the results; detailed per-ablation analyses are provided in Suppl. §F.
@@
Table 7: Ablation study of the synthesis loop (CIFAR-10). 95% confidence intervals are reported in brackets (Wilson score for proportions, $t$-based for means). The full method combines all three components; each ablation removes one.
Removing the novelty filter preserves accuracy trends but roughly halves the number of genuinely distinct architectures admitted (455 $\to$ 220), as the corpus accumulates near-duplicate motifs. Removing the accuracy threshold maintains novelty but weakens the distributional shift: the above-40% fraction drops from 51.1% to 34.0% and mean accuracy falls by 3.8 percentage points, since low-performing models are promoted into the training corpus. Removing iterative fine-tuning entirely (non-iterative baseline) yields the weakest outcome: the valid generation rate, mean accuracy, and above-threshold fraction all remain at the level observed in cycle 1, and only a single novel architecture is discovered because the feedback loop is absent. Taken together, the ablations demonstrate that all three components interact synergistically: the novelty filter sustains exploration, the accuracy threshold steers corpus quality, and iterative fine-tuning closes the feedback loop that enables progressive improvement.
@@ 5 Conclusion and Future Work
This work examines how a code-oriented large language model behaves when placed at the center of an iterative architecture-synthesis loop. Rather than treating the LLM as a fixed component within a neural architecture search pipeline, its output distribution over architectures is tracked across 22 supervised fine-tuning cycles using its own high-quality, structurally novel generations. Under a controlled image-classification setting, the generate–evaluate–select–fine-tune procedure induces a pronounced shift in this distribution, increasing both the likelihood of producing executable models and the early-epoch performance of sampled networks, while retaining non-trivial structural diversity as measured by a code-level novelty filter.
Across cycles, the generator moves from an initial regime in which valid, rapidly learning architectures are relatively uncommon to one in which they constitute the majority of outputs. The distribution of first-epoch accuracies shifts upward: the fraction of models exceeding a moderate performance threshold rises from a small minority early on to a large majority in later cycles. A second outcome is the sustained admission of code-level novel architectures via the MinHash–Jaccard novelty criterion, though we note that text-level novelty does not guarantee functional novelty (see Limitations). A third outcome is the demonstrated generality of these trends: cross-dataset experiments on CIFAR-100 and SVHN confirm that the same qualitative pattern holds across benchmarks varying in number of classes, visual domain, and task difficulty, suggesting that the iterative loop shapes the generator’s code-synthesis priors rather than exploiting dataset-specific artifacts.
Several important questions remain open: whether the proxy ranking under single-epoch evaluation is preserved after full training, whether the discovered architectures are competitive with established NAS methods under matched budgets, and whether the structural diversity documented at the code level (Suppl. §H) extends to a broader set of generated models across cycles. These constitute concrete directions for future work.
@@
Additional future directions include integrating the refined generator with explicit optimization frameworks (e.g., LLMatic, SEKI, or RZ-NAS) to leverage the learned distribution as an architectural prior for downstream search [30, 5, 18]; extending evaluation to higher-resolution datasets and non-classification tasks; incorporating more granular feedback signals such as performance-weighted sampling or reinforcement learning; and integrating multi-objective constraints (e.g., parameter count and latency) to internalize accuracy–efficiency trade-offs critical to edge-oriented NAS [32, 4].